Height of 9-year olds (Independent Samples Z – p – value) At age 9, the average...

  1. Height of 9-year olds (Independent Samples Z – p – value)

At age 9, the average weight (28.6 kg) and height (133.3 cm) for both boys and girls are believed to be exactly the same. A random sample of 9 year-olds yielded these results. Is there a statistically significant difference in the heights of boys and girls at age 9?

Boys

Girls

Sample Size

60

50

Mean height, cm

123.5

126.2

Population variance

98

120

Step 1: State the hypotheses.

Step 2: Compute the test value.

wdOfJw6X8XBjQAAAABJRU5ErkJggg==

Step 3: Find the p – value.

Step 4: Make the decision.

Step 5: Summarize the results.
Answer #1

Step 1) We have to test

Ho : H1-H2 = 0 Ha: Mi-427 0

Step 2) The test statistic is

N V

123.5 - 126.2 z = V 68 + 120 z = -1.3444

Since the averages of boys and girls are assumed to be equal Hi-H2 = 0 . Since the population standard deviations are known, we use z-distribution.

Step 3) Here, Here, the sample means are X1 = 123.5,X2 = 126.2 . The population standard deviations are 6 = 98,0 = 120 . The sample sizes are n = 60, n2 = 50 .

The P-value of the test is

p-value = 2P (Z<z) p-value = 2P(Z < -1.3444) p-value = 0.1788

Step 4) Since p-value = 0.1788 > a = 0.05 , we accept the null hypothesis.

Step 5) We conclude that "We are 95% confident that there is no statistically significant difference in the heights of boys and girls at age 9"